# Parameter estimation for Gibbs distributions

### v9

@@ 4 Solving ${P^{\delta,\varepsilon}_{\tt count}}$ in the continuous setting
1 set $\mathcal{D}\leftarrow{\tt PratioContinuous}(\min\{1/120,\varepsilon/3\},\gamma/4)$.
@@
4       set $\beta_{t}\leftarrow{\tt BalancePreprocessed}(x_{t-1},1/4,\mathcal{D},\frac{\gamma}{10T})$
@@
The complexity bound follows immediately from specification of subroutines. We next analyze the success probability. Throughout, we suppose that that the call at Line 1 succeeds, and produces a good data structure $\mathcal{D}$. There are a number of intermediate calculations.
@@
(a) If $\beta_{t}<\beta_{\max}$, then $\mu_{\beta_{t}}[x_{t-1},n]\geq 1/4$.
(b) If $\beta_{t}>\beta_{\min}$, then $\mu_{\beta_{t}}[0,x_{t-1})\geq 1/4$ and $x_{t}<x_{t-1}$ and $\mu_{\beta_{t}}[0,x_{t})\leq 1/80$ and $\mu_{\beta_{t}}[0,x_{t}]\geq 1/100$.
@@
With probability at least $1-\frac{\gamma}{10T}$, the call to Balance at each iteration $t$ is good. Let us suppose this holds. By definition, this shows (a). For (b), suppose $\beta_{t}>\beta_{\min}$. By specification of Balance we immediately have $\mu_{\beta_{t}}[0,x_{t-1})\geq 1/4$. Let $v,w$ be the minimum values with $\mu_{\beta_{t}}[0,v]\geq 1/80$ and $\mu_{\beta_{t}}[0,w]\geq 1/100$ respectively. (These both exist since the function $x\mapsto\mu_{\beta}[0,x]$ is right-continuous.) Then $\mu_{\beta_{t}}[0,v)\leq 1/80,\mu_{\beta_{t}}[0,w)\leq 1/100$. Since $\mu_{\beta_{t}}[0,x_{t-1})\geq 1/4$, we have $w\leq v<x_{t-1}$.
By Lemma 13, there is a probability of at least $1-\frac{\gamma}{10T}$ that Line 5 well-estimates both intervals $[0,v],[0,w)$ with respect to parameters $\varepsilon/10$ and $\bar{p}=1/100$. Suppose this holds. Then $\hat{\mu}_{\beta_{t}}[0,v]\geq e^{-\varepsilon/10}\mu_{\beta_{t}}[0,v]\geq e^{-1/20}\cdot 1/80>1/90$. So $x_{t}\leq v<x_{t-1}$ and $\mu_{\beta_{t}}[0,x_{t})\leq\mu_{\beta_{t}}[0,v)\leq 1/80$. Likewise, we have $\hat{\mu}_{\beta_{t}}[0,w)\leq e^{1/20}\cdot 1/100<1/90$. Thus $x_{t}\geq w$ and $\mu_{\beta_{t}}[0,x_{t}]\geq\mu_{\beta_{t}}[0,w]\geq 1/100$. ∎
@@
Suppose the algorithm reaches Line 12, and we have $\beta_{1},\dots,\beta_{T}>\beta_{\min}$ with a decreasing sequence $x_{1}>x_{2}>\dots>x_{T-1}>x_{T}$. For each iteration $t<T$, Proposition 23 gives $\mu_{\beta_{t+1}}[x_{t+1},x_{t})=\mu_{\beta_{t+1}}[0,x_{t})-\mu_{\beta_{t+1}}[0,x_{t+1})\geq 1/4-1/80\geq 1/5$ and $\mu_{\beta_{t}}[x_{t+1},x_{t})\leq\mu_{\beta_{t}}[0,x_{t})\leq 1/80$ and $\mu_{\beta_{t+1}}[0,x_{t})\geq 1/4$. Since the map $\beta\mapsto\mu_{\beta}[0,\chi)$ is non-increasing and $\mu_{\beta_{t+1}}[0,x_{t})\geq 1/4>1/80\geq\mu_{\beta_{t}}[0,x_{t})$, we have $\beta_{t+1}<\beta_{t}$ strictly. We can estimate the partition ratio between $\beta_{t+1}$ and $\beta_{t}$ as:
@@
As we show in Lemma 57 in Appendix E, since $b_{1},\dots,b_{g+1}$ is a strictly-increasing positive sequence, we have $\sum_{i=1}^{g}\frac{b_{i}}{b_{i+1}-b_{i}}\geq\frac{g^{2}}{\log(b_{g+1}/b_{1})}-g/2$. Here $b_{g+1}=x_{2}\leq n$, and, since $x_{T-2}<x_{T-1}$ and $x_{T-1}\in\mathcal{F}$, we have $b_{1}\geq 1$. So overall we have shown that
@@
Consider some iteration $t$ and $y\in(x_{t},x_{t-1}]$, where we denote $x_{t}=-\infty$ for $\beta_{t}=\beta_{\min}$. For brevity write $a_{y}=\mu_{\beta_{t}}(y),\hat{a}_{y}=\hat{\mu}_{\beta_{t}}(y)$. By Proposition 23, if $\beta_{t}>\beta_{\min}$ we have $\mu_{\beta_{t}}[0,y)\geq\mu_{\beta_{t}}[0,x_{t}]\geq 1/100$. Likewise, if $\beta_{t}<\beta_{\max}$, we have $\mu_{\beta_{t}}[y,n]\geq\mu_{\beta_{t}}[x_{t-1},n]\geq 1/4$. So $\beta_{t}\in{\tt BalancingVals}(y,1/100)$. By Proposition 10 this shows $a_{y}\geq\Delta(y)/100$. Thus, condition (A2) will hold as long as
@@
For any $y\in\mathcal{F}$, let $\mathcal{E}(y)$ denote the bad event that (11) fails for $y$; note here that we are not necessarily requiring $y\in(x_{t},x_{t-1}]$. Conditioned on $\beta_{t}$, we have
@@ 5 Solving ${P_{\tt count}}$ for integer-valued Gibbs distributions
(b) $k_{1},\dots,k_{t}\in\mathcal{H}$;
(c) $k_{1}<k_{2}<\dots<k_{t}$
(d) $w_{i}\in[0,1]$ for $i=0,\dots,t$.
@@
Our algorithm to solve ${P^{\delta,\varepsilon}_{\tt count}}$ has three stages. As a high-level summary, it proceeds as follows:
  1. 1.
    Construct a suitable covering schedule $\mathcal{I}=(\beta_{0},w_{0},k_{1},\dots,k_{t},\beta_{t},w_{t})$.
  2. 2.
    Estimate the values $Q(\beta_{i})$ for $i=0,\dots,t$.
  3. 3.
    Use these estimates $\hat{Q}(\beta_{i})$ to estimate the counts $c_{i}$.
Using Theorem 6, we can automatically use this algorithm to solve ${P^{{\varepsilon}}_{\tt ratio-all}}$ as well.
The first stage is quite involved, so we defer it to Section 6 where we show the following result:
@@
In the general integer setting, its cost is $O(n(\log^{3}n+\log n\log\tfrac{1}{\gamma}+\log q))$ and ${\tt InvWt}({\cal I})\leq O(n\log n)$.
In the log-concave setting, its cost is $O(n(\log^{2}n+\log\tfrac{1}{\gamma}+\log q))$ and ${\tt InvWt}({\cal I})\leq O(n)$.
For the second stage, we use the following algorithm:
@@
9       for $i=0,\dots,t$ do set $\hat{Q}(\beta_{i})=\hat{Q}(\beta_{i}\mid\mathcal{D})$ for all $i=0,\dots,t$
@@
Note that, somewhat counterintuitively, it can be more efficient to use the continuous-setting algorithm to solve ${P^{{\varepsilon}}_{\tt ratio-all}}$, as opposed to directly estimating the values $\hat{Q}(\beta_{i})$ from the covering schedule itself.
@@
The algorithm ${\tt PratioCoveringSchedule}(\mathcal{I},\varepsilon,\gamma)$ has cost is $O\bigl{(}\frac{\min\{nW,q\log n,q^{2}\}}{\varepsilon^{2}}\log\tfrac{1}{\gamma}\bigr{)}$ for $W={\tt InvWt}(\mathcal{I})$. If $\mathcal{I}$ is proper, then with probability at least $1-\gamma$ it satisfies
@@
Since $t\leq n+1$, Lines $3$ and $4$ have cost $O(nW\varepsilon^{-2}\log\tfrac{1}{\gamma})$. Assuming $\mathcal{I}$ is proper, we have ${{\mathbb{S}}[{X_{i}}]}=\frac{1}{\mu_{\beta_{i-1}}(k_{i})}-1\leq\frac{1}{w_{i-1}}$ for all $i$, so $\sum_{i}{{\mathbb{S}}[{X_{i}}]}\leq W$. Likewise $\sum_{i}{{\mathbb{S}}[{Y_{i}}]}\leq W$. So with probability at least $1-\gamma$, all estimates $\hat{X}^{\tt prod}_{i},\hat{Y}^{\tt prod}_{i}$ are within $e^{\pm\varepsilon/2}$ of $\prod_{j=1}^{i}{\mathbb{E}}[X_{j}],\prod_{j=1}^{i}{\mathbb{E}}[Y_{j}]$ respectively. Observe that
@@ 5.1 Solving ${P^{\delta,\varepsilon}_{\tt count}}$ in the general integer setting
We now move on to the third stage, of using the covering schedule to solve ${P_{\tt count}}$. There are two quite distinct algorithms: one for generic integer-valued distributions, and a specialized algorithm for log-concave distributions. We begin with the algorithm for general integer distributions:
@@
6       find index $i<t$ with $\alpha\in[\beta_{i},\beta_{i+1}]$
@@
Algorithm 8 Solving problem ${P^{\delta,\varepsilon}_{\tt count}}$.
@@
For correctness, assume $\mathcal{I}$ is proper and the calls to Balance and PratioCoveringSchedule are good. Also, assume that Lines 3 and 7 well-estimate every integer in $\mathcal{H}$. By specification of subroutines, these events all hold with probability at least $1-\gamma$. In this case, we claim that the preconditions of Lemma 14 hold for any given $j\in\mathcal{H}$. For this value $j$, let $\alpha,i$ denote the corresponding parameters at Lines 5 and 6 respectively, and define $k=k_{i+1},\beta=\beta_{i}$ for brevity.
By Proposition 10, we have $\mu_{\alpha}(j)\geq\Delta(j)/4$. There are two cases:
@@
  • Suppose $\hat{\mu}_{\alpha}(j)<\delta/4$. Let us assume $j<k$; the case with $j\geq k$ is completely analogous. (A1) holds since $|\log\hat{Q}(\beta)-\log Q(\beta)|\leq\varepsilon/9$. Since Line 7 well-estimates $k$ and $\hat{\mu}_{\alpha}(k)<\delta/4$ we have $\mu_{\alpha}(k)<e^{\varepsilon/9}\cdot\delta/4\leq\delta/2$. We have already seen that $\mu_{\alpha}(j)\geq\Delta(j)/4$ and since $\mathcal{I}$ is proper, we have $\mu_{\beta}(k)\geq w_{i}$. So from the Uncrossing Inequality we get $\mu_{\beta}(j)\geq\frac{\mu_{\beta}(k)\mu_{\alpha}(j)}{\mu_{\alpha}(k)}\geq\frac{w_{i}\Delta(j)/4}{\delta/2}=\frac{w_{i}\Delta(j)}{2\delta}.$ Now Line 7 well-estimates $j$ so

### v11

@@ 4 Solving ${P^{\delta,\varepsilon}_{\tt count}}$ in the continuous setting
1 set $\mathcal{D}\leftarrow{\tt PratioContinuous}(\min\{1/120,\varepsilon/3\},\gamma/4)$
@@
4       set $\beta_{t}\leftarrow{\tt BalancePreprocessed}(x_{t-1},1/4,\mathcal{D},\frac{\gamma}{4T})$
5
@@
The complexity bound follows immediately from specification of subroutines. We next analyze the success probability. Throughout, we suppose that that the call at Line 1 produces a data structure $\mathcal{D}$ which is $\min\{1/120,\varepsilon/3\}$-close. There are a number of intermediate calculations.
@@
(a) If $\beta_{t}<\beta_{\max}$, then $\mu_{\beta_{t}}[x_{t-1},n]\geq 1/4$; if $\beta_{t}>\beta_{\min}$ then $\mu_{\beta_{t}}[0,x_{t-1})\geq 1/4$.
(b) If $\beta_{t}>\beta_{\min}$, then $x_{t}<x_{t-1}$ and $\mu_{\beta_{t}}[0,x_{t})\leq 1/80$ and $\mu_{\beta_{t}}[0,x_{t}]\geq 1/100$.
@@
With probability at least $1-\frac{\gamma}{4T}$, the call to BalancePreprocessed at each iteration $t$ is good. Suppose this holds. By definition, this shows (a). For (b), suppose $\beta_{t}>\beta_{\min}$, and let $v,w$ be the minimum values with $\mu_{\beta_{t}}[0,v]\geq 1/80$ and $\mu_{\beta_{t}}[0,w]\geq 1/100$ respectively. (These both exist since the function $x\mapsto\mu_{\beta}[0,x]$ is right-continuous.) Then $\mu_{\beta_{t}}[0,v)\leq 1/80,\mu_{\beta_{t}}[0,w)\leq 1/100$. Since $\mu_{\beta_{t}}[0,x_{t-1})\geq 1/4$, we have $w\leq v<x_{t-1}$.
By Lemma 13, there is a probability of at least $1-\frac{\gamma}{4T}$ that Line 5 well-estimates both intervals $[0,v],[0,w)$ with respect to parameters $\varepsilon/10$ and $\bar{p}=1/100$. Suppose this also holds. Then $\hat{\mu}_{\beta_{t}}[0,v]\geq e^{-\varepsilon/10}\mu_{\beta_{t}}[0,v]\geq e^{-1/20}\cdot 1/80>1/90$. So $x_{t}\leq v<x_{t-1}$ and $\mu_{\beta_{t}}[0,x_{t})\leq\mu_{\beta_{t}}[0,v)\leq 1/80$. Likewise, we have $\hat{\mu}_{\beta_{t}}[0,w)\leq e^{1/20}\cdot 1/100<1/90$. Thus $x_{t}\geq w$ and $\mu_{\beta_{t}}[0,x_{t}]\geq\mu_{\beta_{t}}[0,w]\geq 1/100$. ∎
@@
Suppose the algorithm reaches Line 12, so $\beta_{1},\dots,\beta_{T}>\beta_{\min}$ and we have a decreasing sequence $x_{1}>x_{2}>\dots>x_{T-1}>x_{T}$. Note that $x_{i}\in\mathcal{F}$ for each $i$, and so $x_{T-1}\geq 1$.
For each iteration $t<T$, Proposition 23 gives $\mu_{\beta_{t+1}}[x_{t+1},x_{t})=\mu_{\beta_{t+1}}[0,x_{t})-\mu_{\beta_{t+1}}[0,x_{t+1})\geq 1/4-1/80\geq 1/5$ and $\mu_{\beta_{t}}[x_{t+1},x_{t})\leq\mu_{\beta_{t}}[0,x_{t})\leq 1/80$ and $\mu_{\beta_{t+1}}[0,x_{t})\geq 1/4$. Since the map $\beta\mapsto\mu_{\beta}[0,\chi)$ is non-increasing and $\mu_{\beta_{t+1}}[0,x_{t})\geq 1/4>1/80\geq\mu_{\beta_{t}}[0,x_{t})$, we have $\beta_{t+1}<\beta_{t}$ strictly. We can estimate the partition ratio between $\beta_{t+1}$ and $\beta_{t}$ as:
@@
As we show in Lemma 58 in Appendix E, since $b_{1},\dots,b_{g+1}$ is a strictly-increasing positive sequence, we have $\sum_{i=1}^{g}\frac{b_{i}}{b_{i+1}-b_{i}}\geq\frac{g^{2}}{\log(b_{g+1}/b_{1})}-g/2$. Here $b_{g+1}=x_{2}\leq n$, and $b_{1}=x_{T-2}\geq 1$. So overall we have shown that
@@
Consider some iteration $t$ and $y\in(x_{t},x_{t-1}]$, where we denote $x_{t}=-\infty$ for $\beta_{t}=\beta_{\min}$. For brevity write $a_{y}=\mu_{\beta_{t}}(y),\hat{a}_{y}=\hat{\mu}_{\beta_{t}}(y)$. By Proposition 23, if $\beta_{t}>\beta_{\min}$ we have $\mu_{\beta_{t}}[0,y)\geq\mu_{\beta_{t}}[0,x_{t})\geq 1/100$. Likewise, if $\beta_{t}<\beta_{\max}$, we have $\mu_{\beta_{t}}[y,n]\geq\mu_{\beta_{t}}[x_{t-1},n]\geq 1/4$. So $\beta_{t}\in{\tt BalancingVals}(y,1/100)$. By Proposition 12 this shows $a_{y}\geq\Delta(y)/100$. Thus, condition (A2) will hold as long as
@@
For any $y\in\mathcal{F}$, let $\mathcal{E}_{y}$ denote the bad event that (11) fails for $y$; note here that we are not necessarily requiring $y\in(x_{t},x_{t-1}]$. Conditioned on $\beta_{t}$, we have
@@ 5 Solving ${P_{\tt count}}$ for integer-valued Gibbs distributions
(b) $k_{1}<k_{2}<\dots<k_{t}$ with $k_{1},\dots,k_{t}\in\mathcal{H}$;
(c) $w_{i}\in[0,1]$ for $i=0,\dots,t$.
@@
It is quite involved to actually generate such a covering schedule, so we defer the technical details to Section 6 where we show the following:
@@
In general integer setting, it has cost $O(n(\log^{3}n+\log n\log\tfrac{1}{\gamma}+\log q))$ and gives ${\tt InvWt}({\cal I})\leq O(n\log n)$.
In the log-concave setting, it has cost $O(n(\log^{2}n+\log\tfrac{1}{\gamma}+\log q))$ and gives ${\tt InvWt}({\cal I})\leq O(n)$.
For the moment, let us suppose we are given such a covering schedule. There are then two main steps to solve ${P^{\delta,\varepsilon}_{\tt count}}$. First, we use a telescoping-product calculation similar to the Paired Product Estimator to estimate the values $Q(\beta_{i})$ for $i=0,\dots,t$. Next, we use these estimates $\hat{Q}(\beta_{i})$ to interpolate the counts $c_{i}$.
Via Theorem 6, one may use the solution for ${P^{\delta,\varepsilon}_{\tt count}}$ with $\delta=1/n$ to also solve ${P_{\tt ratio-all}}$.
We now describe these steps in detail. First, we develop the algorithm PratioCoveringSchedule to estimate the values $Q(\beta_{i})$ in the covering schedule.
@@
9       for $i=0,\dots,t$ do set $\hat{Q}(\beta_{i})=\hat{Q}(\beta_{i}\mid\mathcal{D})$
@@
The algorithm ${\tt PratioCoveringSchedule}(\mathcal{I},\varepsilon,\gamma)$ has cost $O\bigl{(}\frac{\min\{nW,q\log n,q^{2}\}}{\varepsilon^{2}}\log\tfrac{1}{\gamma}\bigr{)}$ for $W={\tt InvWt}(\mathcal{I})$. If $\mathcal{I}$ is proper, then with probability at least $1-\gamma$ it satisfies
@@
Since $t\leq n+1$, Lines $3$ and $4$ have cost $O(\frac{nW}{\varepsilon^{2}}\log\tfrac{1}{\gamma})$. Assuming $\mathcal{I}$ is proper, we have ${{\mathbb{S}}[{X_{i}}]}=\frac{1}{\mu_{\beta_{i-1}}(k_{i})}-1\leq\frac{1}{w_{i-1}}$ for all $i$, so $\sum_{i}{{\mathbb{S}}[{X_{i}}]}\leq W$. Likewise $\sum_{i}{{\mathbb{S}}[{Y_{i}}]}\leq W$. So with probability at least $1-\gamma$, all estimates $\hat{X}^{\tt prod}_{i},\hat{Y}^{\tt prod}_{i}$ are within $e^{\pm\varepsilon/2}$ of $\prod_{j=1}^{i}{\mathbb{E}}[X_{j}],\prod_{j=1}^{i}{\mathbb{E}}[Y_{j}]$ respectively. Observe that
@@
We next turn to using these estimates to finish solving the problem ${P^{\delta,\varepsilon}_{\tt count}}$. We develop two separate algorithms: one for the general integer setting, and a second optimized for the log-concave setting.
@@ 5.1 Solving ${P^{\delta,\varepsilon}_{\tt count}}$ in the general integer setting
6       find index $i\in\{0,\dots,t-1\}$ with $\alpha\in[\beta_{i},\beta_{i+1}]$
@@
Algorithm 8 Solving problem ${P^{\delta,\varepsilon}_{\tt count}}$ for the general integer setting.
@@
For correctness, assume $\mathcal{I}$ is proper and the calls to Balance and PratioCoveringSchedule are good. Also, assume that Lines 3 and 7 well-estimate every value in $\mathcal{H}$. By specification of subroutines, these events all hold with probability at least $1-\gamma$. In this case, we claim that the preconditions of Lemma 14 hold for all $j\in\mathcal{H}$. To show this, consider $j\in\mathcal{H}$, and let $\alpha,i$ denote the corresponding parameters at Lines 5 and 6 respectively, and define $k=k_{i+1},\beta=\beta_{i}$ for brevity. By Proposition 12, we have $\mu_{\alpha}(j)\geq\Delta(j)/4$. There are two cases:
@@
  • Suppose $\hat{\mu}_{\alpha}(j)<\delta/4$. Let us assume $j<k$; the case with $j\geq k$ is completely analogous. (A1) holds since $|\log\hat{Q}(\beta)-\log Q(\beta)|\leq\varepsilon/9$. Since Line 7 well-estimates $k$ and $\hat{\mu}_{\alpha}(k)<\delta/4$ we have $\mu_{\alpha}(k)<e^{\varepsilon/9}\delta/4\leq\delta/2$. We have already seen that $\mu_{\alpha}(j)\geq\Delta(j)/4$ and since $\mathcal{I}$ is proper, we have $\mu_{\beta}(k)\geq w_{i}$. So from the Uncrossing Inequality we get $\mu_{\beta}(j)\geq\frac{\mu_{\beta}(k)\mu_{\alpha}(j)}{\mu_{\alpha}(k)}\geq\frac{w_{i}\Delta(j)/4}{\delta/2}=\frac{w_{i}\Delta(j)}{2\delta}.$ Now Line 7 well-estimates $j$ so